# MATTER: Multiscale Attention for Registration Error Regression

### v2

@@ 1 Introduction
Point cloud registration (PCR) is a fundamental task in three-dimensional (3D) computer vision, aiming to estimate the rigid transform between two overlapping point clouds. It serves as a backbone for numerous downstream applications, including simultaneous localization and mapping (SLAM), 3D reconstruction, and robotic navigation. Despite recent advances in PCR, registration error is still significant in challenging scenarios, e.g., when optimization of the loss function gets trapped in local minima [1], under motion distortion [2] or measurement noise, or when the problem is geometrically under-constrained [3, 4]. Errors in PCR can propagate to downstream modules, resulting in map distortion and navigation failure. This motivates the need for registration misalignment detection, i.e., predicting, after registration, whether the registered point clouds deviate from the well-aligned ground truth. This prediction can then allow downstream tasks to take corrective action.
@@
The objective functions used in existing registration methods, such as closest-point residual [5], score function [6], and entropy gain [7], do not reflect the alignment error, as they are not evaluated using the ground truth pose. Recent works [8, 9, 10, 11, 12, 13] proposed learning-based approaches for point cloud misalignment classification (PCMC), where the registration result is categorized into discrete quality levels. Specifically, in [12, 8], logistic regression is used with manually designed features for binary classification. CorAl, introduced in [13], fits the neighborhood of each point to a 3D Gaussian distribution and uses the differential entropy of this distribution as the feature. FACT [11]—the state-of-the-art method—extends binary classification to multiple classes, which allows evaluation at a finer level. However, all these works are limited to classification, i.e., coarsely categorizing the misalignment into a few levels, without providing continuous predictions of the actual alignment error.
In this paper, we cast registration error estimation as a continuous prediction problem, regressing a scalar alignment error from multiscale geometric features fused by attention. Our contributions are threefold. First, we extend PCMC to point cloud misalignment regression (PCMR) for fine-grained assessment. Second, we propose multiscale attention for registration error regression (MATTER), which extracts features at multiple geometric scales and fuses them using attention. We depict the proposed attention architecture in Fig. 1, and detail it in Section 2. This significantly improves robustness over single-scale features. Finally, we adapt prior PCMC methods to regression and show, via numerical experiments, that MATTER consistently outperforms them on three datasets. We also consider a mapping downstream task and use MATTER to detect misaligned frames that need correction, i.e., to be re-registered. Notably, MATTER achieves a better map quality than the state-of-the-art classification approach for the same amount of re-registration.
@@ 2.2 Model Architecture
We adapt the model architecture in FACT [11] to our PCMR task by using multiscale attention to fuse alignment-related features. See the illustration in Fig. 1. First, we follow [11] to extract global features and local features in local neighborhoods at multiple scales. Afterwards, both types of features are concatenated and cross-attended to extract scale-wise attention score. The attention score is then used as a weighting factor for multiscale features before passing into a point transformer and regression head for the final prediction. Our model is then trained with an $l_{2}$ loss.
@@ 2.3 Feature Extraction
We first transform both $P^{\mathrm{S}}$ and $P^{\mathrm{R}}$ point clouds into a common frame and apply the farthest point sampling [14] to select the same number of anchor points from each of them. In the rest of the section, we focus on a generic anchor point. Let $\mathcal{N}_{P^{\mathrm{S}}}$ and $\mathcal{N}_{P^{\mathrm{R}}}$ denote the points from the source and reference point clouds, respectively, within the neighborhood of this anchor point. We use $\mathcal{N}_{\mathrm{sep}}$ to refer to $\mathcal{N}_{P^{\mathrm{S}}}$ if the anchor point belong to $P^{\mathrm{S}}$, and to $\mathcal{N}_{P^{\mathrm{R}}}$ if it belongs to $P^{\mathrm{R}}$. Define the joint neighborhood $\mathcal{N}_{\mathrm{joint}}=\mathcal{N}_{P^{\mathrm{S}}}\cup\mathcal{N}_{P^{\mathrm{R}}}$. Following [11], we extract local features around each anchor, namely the separate and joint differential entropy [13] $H(\mathcal{N}{\mathrm{sep}}),H(\mathcal{N}{\mathrm{joint}})$, and the Sinkhorn divergence $D_{\lambda}$. They measure the discrepancy between the distributions of points in the aligned source with respect to reference neighborhoods. Additional details on the feature computation are provided in Appendix. A. Unreliable neighborhoods are handled explicitly: if one point cloud contributes no points, the differential entropy and Sinkhorn divergence are set to a bounded default. The next local features are coverage ratios $\rho_{\mathrm{joint}}=\big|\mathcal{N}_{\mathrm{joint}}\big|/(\big|P^{\mathrm{S}}\big|+\big|P^{\mathrm{R}}\big|)$ and $\rho_{\mathrm{sep}}=\big|\mathcal{N}_{\mathrm{sep}}\big|/\big|P_{\mathrm{sep}}\big|$, where $\big|P_{\mathrm{sep}}\big|$ is given by $\big|P^{\mathrm{{S}}}\big|$ or $\big|P^{\mathrm{{R}}}\big|$ if the anchor point belongs to $P^{\mathrm{{S}}}$ or $P^{\mathrm{{R}}}$, respectively.
Global features that are computed on the whole point cloud include a co-visibility score $c\!\in[0,1]$ computed with a visibility operator [15], the distance $d$ from the point to the LiDAR sensor, and a binary source flag $b\in\{0,1\}$ designating which point cloud the point originates from. These features serve to down-weight anchor points with non-covisible or distant points, while up-weighting well-observed ones.
@@ 3 Evaluation
To approximate real-world deployment, we construct realistic datasets and compare against state-of-the-art misalignment classifiers. For the regression task, we adapt these classifiers by replacing their classification heads with a scalar predictor and training them end-to-end with an $\ell_{2}$ loss on our labels.
@@ 3.2 Evaluation Results
We train and evaluate on the three datasets described above and report the root mean square error (RMSE), the mean absolute error (MAE), and the coefficient of determination ($R^{2}$). For our method, we consider three radii $r_{1}=7.5$ $\mathrm{m},$ $r_{2}=4.0\,\mathrm{m},$ $r_{3}=2.5\,\mathrm{m}$, and use temperature $\tau=0.6$. We report the results of our method with different $\tau$ in Appendix. B.
For all baselines, we follow the same settings as reported in the original papers, e.g., FACT and CorAl both use a single adaptive radius within $[0.5,7.5]$ m. Results in Table 1 show that our method achieves the best performance on all datasets in terms of all three metrics RMSE, MAE, and $R^{2}$. Notably, on nuScenes-ICP, most methods have $\mathrm{RMSE}$ slightly above $\mathrm{MAE}$, indicating few large outliers. In contrast, on KITTI–GeoTransformer, $\mathrm{RMSE}$ is markedly larger than $\mathrm{MAE}$, indicating heavier-tailed errors that are harder to detect.
To better understand the cause of the improvement, we visualize the the registered point clouds in Fig. 2, and the per-anchor-point scale that is most highly weighted by MATTER’s attention. MATTER adaptively favors different radii, while FACT sets its single radius to $7.5$ m for all anchor points in Fig. 2(2(a)) and $95\%$ of the anchor points in Fig. 2(2(b)). Specifically, our attention mechanism favors the smallest scale S1 in dense, well-overlapping regions where the local geometry better approximates a 3D Gaussian, making the closed-form differential-entropy descriptors more reliable. When correspondences are absent or alignment is poor, the model adaptively enlarges the radius. In Fig. 2(2(a)), the left side of the reference point cloud lacks corresponding points from the source, which triggers the largest scale S3 so that anchors still cover areas that contain matches. In Fig. 2(2(b)), the largest scale S3 is favored in the red circles because small neighborhoods fail to capture counterparts of both clouds. The midscale S2 is generally less preferred but appears in cluttered medium-thickness structures.
@@
Fig. 3: Visualization of re-registered point cloud sequences. Left: FACT ($E_{\text{align}}=21.46$ m); Right: MATTER ($E_{\text{align}}=18.35$ m).
@@ 3.3 Downstream Task
Beyond comparison with a fixed threshold, we plot in Fig. 4 the final-frame alignment error under various re-registration rates (similar to [21, Fig. 5]). By correcting more frames, the final-frame alignment error keeps reducing. For low re-registration rates (e.g., from 7% to 40%), MATTER reaches significantly lower alignment error of the final frame by correcting the same amount of previous frames.
@@ 4 Conclusion
Appendix A Feature Extraction
@@
Appendix B Ablation on Temperature

### v3

@@ 1 Introduction
Point cloud registration (PCR) is a core problem in 3D computer vision, aiming to estimate the rigid transform between two overlapping point clouds. It underpins numerous downstream applications, including simultaneous localization and mapping (SLAM), 3D reconstruction, and robotic navigation. Despite recent advances in PCR, registration error is still significant in challenging scenarios, e.g., when optimization of the loss function gets trapped in local minima [17], under motion distortion [20] or measurement noise, or when the problem is geometrically under-constrained [15, 3]. Errors in PCR can propagate to downstream modules, resulting in map distortion and navigation failure. This motivates the need for registration misalignment detection, i.e., predicting, after registration, whether the registered point clouds deviate from the well-aligned ground truth. This prediction can then allow downstream tasks to take corrective action.
@@
The objective functions used in existing registration methods, such as closest-point residuals [7], score functions [13], and entropy gain [18], are not evaluated against the ground truth pose and are sensitive to local minima [17], noise, and scene geometry, their final values do not reliably reflect the true alignment quality. Recent works [2, 4, 6, 8, 19, 1] proposed learning-based approaches for point cloud misalignment classification (PCMC), where the registration result is categorized into discrete quality levels. Specifically, in [19, 2], logistic regression is used with manually designed features for binary classification. CorAl, introduced in [1], fits a 3D Gaussian distribution to the neighborhood of each point, and uses the differential entropy of this distribution as the feature. FACT [8]—the state-of-the-art method—extends binary classification to multiple classes, which allows evaluation at a finer level. However, all these works are limited to classification, i.e., coarsely categorizing the misalignment into a few levels, without providing continuous predictions of the actual alignment error. A more comprehensive review of related work is provided in Appendix B.
We here recast PCMC as a continuous prediction problem, regressing a scalar alignment error from multiscale geometric features fused by attention. Our contributions are threefold. First, we extend PCMC to point cloud misalignment regression (PCMR) for fine-grained assessment. Second, we introduce MATTER (Multiscale ATTention for registration Error Regression), which improves robustness by extracting features at multiple geometric scales and fusing them using attention. The MATTER architecture is shown in Fig. 1, and detailed in Section 2. Third, we adapt prior PCMC methods to regression and demonstrate through experiments on three datasets, that MATTER consistently outperforms them. We also apply MATTER to a mapping downstream task, to detect misaligned frames that need correction. Under the same re-registration budget, MATTER achieves a better map quality than the state-of-the-art methods.
@@ 2.2 Model Architecture
We adapt the FACT architecture [8] to the PCMR task by using multiscale attention to fuse alignment-related features, see Fig. 1. Following [8], we extract global and local neighbourhood features at multiple scales. These features are then concatenated and cross-attended to extract scale-wise attention weights. The weights are then used to reweight the multiscale features before passing to a point transformer and regression head for the final prediction. Training uses an $l_{2}$ loss.
@@ 2.3 Feature Extraction
We first transform both $P^{\mathrm{S}}$ and $P^{\mathrm{R}}$ point clouds into a common frame and apply farthest point sampling [9] to select the same number of anchor points from each of them. In the rest of the section, we focus on a generic anchor point. Let $\mathcal{N}_{P^{\mathrm{S}}}$ and $\mathcal{N}_{P^{\mathrm{R}}}$ denote the points from the source and reference point clouds, respectively, within the neighborhood of this anchor point. We use $\mathcal{N}_{\mathrm{sep}}$ to refer to $\mathcal{N}_{P^{\mathrm{S}}}$ if the anchor point belong to $P^{\mathrm{S}}$, and to $\mathcal{N}_{P^{\mathrm{R}}}$ if it belongs to $P^{\mathrm{R}}$. Define the joint neighborhood $\mathcal{N}_{\mathrm{joint}}=\mathcal{N}_{P^{\mathrm{S}}}\cup\mathcal{N}_{P^{\mathrm{R}}}$. Following [8], we extract local features around each anchor, namely the separate and joint differential entropy [1] $H(\mathcal{N}{\mathrm{sep}}),H(\mathcal{N}{\mathrm{joint}})$, and the Sinkhorn divergence $D_{\lambda}$. They measure the discrepancy between the distributions of points in the aligned source with respect to reference neighborhoods. Additional details on the feature computation are provided in Appendix C. Unreliable neighborhoods are handled explicitly: if one point cloud contributes no points, the differential entropy and Sinkhorn divergence are set to a bounded default. The next local features are coverage ratios $\rho_{\mathrm{joint}}=\big|\mathcal{N}_{\mathrm{joint}}\big|/(\big|P^{\mathrm{S}}\big|+\big|P^{\mathrm{R}}\big|)$ and $\rho_{\mathrm{sep}}=\big|\mathcal{N}_{\mathrm{sep}}\big|/\big|P_{\mathrm{sep}}\big|$, where $\big|P_{\mathrm{sep}}\big|$ is given by $\big|P^{\mathrm{{S}}}\big|$ or $\big|P^{\mathrm{{R}}}\big|$ if the anchor point belongs to $P^{\mathrm{{S}}}$ or $P^{\mathrm{{R}}}$, respectively.
Three global point features are computed on the whole point cloud: (1) a co-visibility score $c\!\in[0,1]$ computed with a visibility operator [11], (2) the distance $d$ from the point to the LiDAR sensor, and (3) a binary source flag $b\in\{0,1\}$ designating which point cloud the point originates from. These features serve to down-weight non-covisible or distant regions, while up-weighting well-observed ones.
@@ 3 Evaluation
To approximate real-world deployment, we construct realistic datasets and compare against state-of-the-art misalignment classifiers. For the regression task, we adapt these classifiers by replacing their classification heads with a scalar predictor and training them end-to-end with an $\ell_{2}$ loss on our labels. The training details are provided in Appendix A.
@@ 3.2 Evaluation Results
We train and evaluate on the three datasets described above and report the root mean square error (RMSE), the mean absolute error (MAE), and the coefficient of determination ($R^{2}$). For our method, we consider three radii $r_{1}=7.5$ $\mathrm{m},$ $r_{2}=4.0\,\mathrm{m},$ $r_{3}=2.5\,\mathrm{m}$, and use temperature $\tau=0.6$. We report the results of our method with different $\tau$ in Appendix. D.
For all baselines, we follow the same settings as reported in the original papers. e.g., FACT and CorAl both use a single adaptive radius within $[0.5,7.5]$ m. Results in Table 1 show that our method achieves the best performance on all datasets in terms of all three metrics RMSE, MAE, and $R^{2}$. Notably, on nuScenes-ICP, most methods have $\mathrm{RMSE}$ slightly above $\mathrm{MAE}$, indicating few large outliers. In contrast, on KITTI–GeoTransformer, $\mathrm{RMSE}$ is markedly larger than $\mathrm{MAE}$, indicating heavier-tailed errors that are harder to detect.
To better understand the cause of the improvement, we visualize the registered point clouds in Fig. 2, and the per-anchor-point scale that is most highly weighted by MATTER’s attention. MATTER adaptively favors different radii, while FACT sets its single radius to $7.5$ m for all anchor points in Fig. 2(2(a)) and $95\%$ of the anchor points in Fig. 2(2(b)). Specifically, our attention mechanism favors the smallest scale S1 in dense, well-overlapping regions where the local geometry better approximates a 3D Gaussian, making the closed-form differential-entropy descriptors more reliable. When correspondences are absent or alignment is poor, the model adaptively enlarges the radius. In Fig. 2(2(a)), the left side of the reference point cloud lacks corresponding points from the source, which triggers the largest scale S3 so that anchors still cover areas that contain matches. In Fig. 2(2(b)), the largest scale S3 is favored in the red circles because small neighborhoods fail to capture counterparts of both clouds. The midscale S2 is generally less preferred but appears in cluttered medium-thickness structures.
@@
Fig. 3: Visualization of re-registered point cloud sequences. Left: FACT, $E_{\text{align}}=21.46$ m; Right: MATTER, $E_{\text{align}}=18.35$ m.
@@ 3.3 Downstream Task
Beyond comparison with a fixed threshold, we plot in Fig. 4 the final-frame alignment error under various re-registration rates (similar to [12, Fig. 5]). By correcting more frames, the final-frame alignment error keeps reducing. For low re-registration rates (e.g., from 7% to 40%), MATTER reaches significantly lower alignment error of the final frame by correcting the same amount of previous frames.
@@ 4 Conclusion
Appendix A Training Settings
All models were trained on an NVIDIA A100 GPU using the Adam optimizer with a batch size of 16 for 200 epochs. The initial learning rate was set to $1\times 10^{-4}$ with weight decay $1\times 10^{-4}$. A step learning-rate scheduler was employed, decaying the learning rate every 10 epochs with a decay factor of 0.825. The backbone network followed the PointTransformer [21] architecture consisting of 4 transformer–transition blocks and a transformer feature dimension of 256. During training, data augmentation included random dropout, scaling, and translation. For point sampling, farthest point sampling (FPS) was applied to obtain 2048 points per point cloud. Differential entropy estimation used a logarithmic stability parameter $\log\epsilon=-18$ with a rejection threshold of 0.20. Sinkhorn divergence was computed with $p=2$ and blur parameter 0.05.
Appendix B Related Works
Existing approaches for post-registration quality assessment can be broadly divided into objective-function-based and learning-based methods. Classical registration pipelines typically rely on the final value of the objective during optimization, such as closest-point residuals [7], score functions [13], or entropy gain [18], as an indicator of alignment quality. However, these objectives are designed to measure fitness during optimization rather than to measure the true pose error, and their values are highly sensitive to local minima, noise, and scene geometry. As a result, they do not exhibit a reliable or monotonic relationship with the actual alignment error [17]. As noted in [1], such metrics “provide limited information on whether the point clouds are correctly aligned once registration has been carried out,” meaning that a low objective value does not necessarily imply correct alignment.
To overcome this limitation, several works have proposed learning-based point cloud misalignment classification (PCMC) methods, which predict the registration quality directly from geometric features. Early approaches [19, 2] employ manually designed features combined with logistic regression to perform binary classification of alignment quality. CorAl [1] improves feature representation by fitting a local 3D Gaussian distribution to each point neighborhood and using the resulting differential entropy as a descriptor. More recently, FACT [8] extends the formulation from binary to multi-class classification, enabling a finer categorization of misalignment levels and representing the current state of the art in PCMC.
Despite these advances, existing PCMC methods remain limited to discrete classification, providing only coarse alignment quality labels rather than continuous estimates of the true registration error. This motivates the need for point cloud misalignment regression, which directly predicts the alignment error in a continuous domain.
Appendix C Feature Extraction
@@
Appendix D Ablation on Temperature
@@ Appendix E Larger Visualizations
For improved readability and interpretation, we include larger versions of Fig. 2 and Fig. 3 from the main paper in Fig. 5.
(a) Poor registration on KITTI with $E_{\text{align}}=0.08$ m. The bottom-left area is non-overlapping. Colored circles visualize neighborhood ranges of the corresponding scales.
(b) Poor registration on nuScenes with $E_{\text{align}}=2.46$ m. The red circle highlights a challenging region where larger neighborhoods are preferred.
Fig. 5: Larger visualizations from the main paper. Top: registered point clouds and scale preference. Bottom: re-registered point cloud sequences (Left: FACT, $E_{\text{align}}=21.46$ m; Right: MATTER, $E_{\text{align}}=18.35$ m).